# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v2

@@ Abstract
In standard treatments of electron transport, momentum relaxation in a perfect, defect-free crystal is linked with phonon creation or annihilation. In this work, we reconsider this problem for a finite, isolated crystal, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum together with discrete crystal pseudomomentum. Starting from the density–density form of the electron–lattice interaction, we show that an electron in the interior of a perfect crystal admits elastic momentum-transfer channels in which total momentum is conserved by recoil of the lattice background without phonon excitation. These elastic channels can provide the leading contribution to momentum relaxation. We further identify mixed quasielastic and superelastic processes in which phonon occupations change but do not account entirely for the electron’s momentum transfer. The elastic channels arise within the standard microscopic Hamiltonian and do not require additional disorder or defects. The resulting framework provides a complementary microscopic perspective on momentum relaxation in clean crystals, and is consistent with experimental phenomena such as weak localization, quantum oscillations, ultrasonic attenuation, and the observed separation of momentum and energy relaxation times.
@@ I Introduction
Electron–lattice interactions play a central role in condensed matter physics. They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence. Experimentally, however, momentum and energy relaxation often occur on markedly different timescales. In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow. The resulting hierarchy
@@
In conventional transport theory, momentum relaxation in a perfect, defect-free crystal is typically described in terms of phonon-mediated scattering processes. Within the Bloch–Grüneisen and semiclassical Boltzmann frameworks, transport is treated as a sequence of transitions between electronic states, with rates evaluated using the electron–phonon interaction [35, 3]. In the standard Fröhlich Hamiltonian [9], the interaction is expressed explicitly in terms of phonon creation and annihilation operators, and momentum transfer is therefore associated, at lowest order, with changes in phonon occupation. In this commonly employed description, elastic momentum relaxation is primarily attributed to impurities or other forms of explicit disorder.
The observed separation between $\tau_{p}$ and $\tau_{E}$ motivates a careful examination of how momentum conservation is implemented microscopically in a perfectly periodic crystal. The conceptual distinction between momentum relaxation and energy dissipation has long been recognized. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted that the lattice, as a massive collective object, can absorb momentum with minimal energy cost [24]. Pippard similarly stressed that while defects often play an important role, momentum relaxation need not be uniquely tied to strongly inelastic processes [25].
In this work we revisit the problem starting from the same microscopic electron–lattice Hamiltonian, but retaining explicitly the lattice center-of-mass degree of freedom. In conventional treatments the crystal is effectively considered as a fixed reference frame, and the translational zero mode does not appear among the dynamical variables. Here we adopt an expanded viewpoint in which the crystal total mechanical momentum is retained. This does not modify the underlying Hamiltonian; rather, it makes explicit a degree of freedom that is commonly suppressed.
When total momentum conservation is enforced in this manner, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This exchange can occur without excitation of internal phonon modes, leading to an elastic recoil channel even in a perfectly periodic crystal. We also identify mixed quasielastic and superelastic processes in which phonon occupations change but do not fully account for the electron’s momentum transfer. These channels arise within the standard density–density form of the electron–lattice interaction [22, 9] and require no additional disorder.
The physical content of the elastic recoil channel is elementary. In any untethered composite object, an impulse imparted to one constituent immediately changes the total momentum of the whole. While redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the collective momentum of the backgroud crystal appears in the wavefunction and is not dependent of the details of redistribution of the momentum among the constitutuents of the whole. . An instructive analogue is provided by Mössbauer scattering, in which the dominant elastic line is recoil of the entire lattice without internal phonon excitation.
A conceptual roadblock has been that, in contrast to processes such as Mössbauer and neutron scattering, where elastic, non-Bragg diffuse deflection is widely acknowledged, conduction electrons do not enter and exit; they start and stay in the crystal. The roadblock is addressed by assigning pseudomomentum to both a “foreground” electron density and a “background” consisting of everything else. We hasten to point out that the question here involves real physics, and not mere representation: Must phonons be created or destroyed when electrons deflect in a pure crystal, or not?
On the meaning of elastic scattering
Elastic scattering of interior electrons is most familiar in the context of impurity or defect scattering, where translational symmetry is explicitly broken. Here, we restore pseudomomentum to the status of a conserved momentum, and as in free space what then is conserved in such an elastic deflection? The answer will not be controversial: the total momentum of the whole system is conserved. What we believe is new is that two pseudmomomenta have changed equally and oppositely in such a collision: one belonging to the electron, another to the rest of the crystal. The total momentum, the sum of the two, is unchanged. In a perfectly periodic crystal, the relationship between changes in electronic pseudomomentum and the mechanical momentum of the lattice background is less often made explicit. In the present formulation this bookkeeping becomes transparent once (i) the lattice is treated as fully dynamical, (ii) pseudomomentum is assigned both to the electron and to the background in the absence of the electron density, and (iii) the interaction is expressed in density–density form, making the exchange of pseudomomentum explicit.

### v3

# Elastic and Quasielastic Electron Scattering in Perfect Crystals

@@ Abstract
Momentum relaxation of electrons in perfect crystals is conventionally associated with phonon emission or absorption and is therefore regarded as intrinsically inelastic. We show that this is not a fundamental requirement. Starting from the Fröhlich Hamiltonian, but making fewer assumptions than usual—retaining total mechanical momentum and the full density-density electron-lattice interaction—we find a substantially larger class of allowed processes. When total momentum conservation is enforced together with all pseudomomenta, the lattice center-of-mass provides a continuous momentum reservoir, and phonons are no longer required to account for the electron’s momentum change. The conventional phonon-mediated processes emerge as a restricted subset of a higher-dimensional manifold of transitions. Within this expanded space, elastic and quasielastic momentum-transfer channels appear already at first order, even in a perfect crystal. Momentum relaxation can therefore proceed rapidly while energy relaxation remains slow, consistent with a wide range of experimental observations including hot-electron injection, quantum oscillations, weak localization, and Coulomb drag in low-dimensional conductors. In particular, scattering in a perfect one-dimensional wire is not kinematically constrained once the lattice recoil degree of freedom is retained.
@@ I Introduction
Electron–lattice interactions play a central role in condensed matter physics. They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence.
In the modern transport theory of metals (Bloch–Peierls–Ziman–Mott–Ashcroft–Mermin), the electron–lattice problem is formulated by imposing periodic boundary conditions and expanding lattice displacements using internal normal modes. In this construction, going back to Born and von K’arm’an, the lattice center-of-mass (zero) mode is removed at the outset, leaving only internal vibrational degrees of freedom. When an electron applies force to a lattice atom, it can only respond with finite energy phonons. The zero energy recoil, without which Mössbauer scattering would falter, has been removed. In first order, momentum transfer is forced to reside in a single phonon, leading to the familiar selection rules $k’=k\pm q$ and $\varepsilon_{k’}=\varepsilon_{k}\pm\hbar\omega_{q}$. In a perfect crystal, a change in electron momentum thus requires phonon emission or absorption and is therefore inelastic. Although this is true within the formulation, it is an unnatural result, not followed in the laboratory.
When the lattice center-of-mass degree of freedom is retained, the total pseudomomentum is conserved and can be exchanged between the electron and the background. If the zero mode is removed, this exchange channel is absent. The phonon sector alone does not form a closed system for pseudomomentum, and the familiar selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm$
need not hold for individual processes. Indeed, it is violated once the density-density interaction is treated beyond the linear Fröhlich approximation. Pseudomomentum is therefore not a conserved quantity within the electron-phonon subspace alone. The traditional clean pseudomomentum accounting falters if even the next terms are kept in the the density-density expansion, e.g.
Purely elastic processes are a rarity in the classical world, but are commonplace in the quantum world. Consider a linear atom-diatom collision, i.e. three translational degrees of freedom, separating to a center mass of the diatom, the atom, and the internal virbatipnal coordinate: if not too violent, upon measurement the diatom is likely to be in the state it started in, even though it has definitely been disturbed. Again, Mössbauer emission comes to mind. In fact for pass-through experiments, like fast electrons, both Bragg and a good portion of diffuse scattering is known to be elastic, with the crystal recoiling as a whole. Somehow an electron residing inside a crystal has found itself treated a different way, but it should not be.
Experimentally, momentum and energy relaxation often occur on markedly different timescales. In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow. The resulting hierarchy
@@
The experimentally observed separation between $\tau_{p}$ and $\tau_{E}$ does not seem to comport with the “always inelastic” narrative, and this reality check motivates a careful examination of how momentum conservation is implemented in a perfectly periodic crystal.
At finite temperature, lattice vibrations introduce perturbations that allow electrons to scatter between Bloch states. This mechanism was first analyzed by Bloch, who obtained the characteristic low-temperature $T^{5}$ resistivity associated with electron-phonon scattering in metals [8]. These ideas were incorporated into the Bloch-Grüneisen-Boltzmann transport framework [8, 13], in which the electronic distribution function evolves according to a Boltzmann equation for Bloch electrons. Within this formulation, scattering is described as transitions between Bloch states accompanied by phonon absorption or emission, enforced by energy-conserving delta functions. In the standard Fröhlich Hamiltonian [10], the interaction is expressed in terms of phonon creation and annihilation operators, so that momentum transfer is associated, at lowest order, with changes in phonon occupation and is therefore intrinsically inelastic, subject to pseudomomentum and energy conservation [39, 4]. Elastic momentum relaxation is accordingly attributed to impurities or other forms of explicit disorder. The subsequent development of Migdal-Eliashberg theory further reinforced this picture, embedding the identification of momentum relaxation with inelastic electron-phonon scattering into the standard theoretical framework.
The situation has been different for external, “pass through” probes such as X-rays or neutrons or even fast electrons. In such circumstances, elastic and even diffuse elastic recoil of the entire crystal is routinely included through Debye-Waller theory. Here, we will see that there is no solid basis for the asymmetry in the treatment of an interior electron.
As we will see, bringing back the center of mass dramatically changes the narrative of internal scattering processes. Somehow that seems implausible, which is one reason perhaps that the inelastic narrative has persisted for so long.
We revisit this issue determined to be rigorous about momentum conservation in a fully isolated electron-crystal system with all applicable momenta retained. (The isolation can later be broken once a full understanding is reached of the isolated system.) When the electron-lattice interaction is written in its full density-density form, elastic and quasi-elastic momentum exchange channels appear already at first order, without any defects present other than the transient ones implied by the thermal deformation potential.
A few preliminary, qualitative remarks are in order. First, removing the center of mass makes electrons in crystals is akin to the venerable particle in a box model. There too there is no center of mass of the box, and no box recoil. Momentum is not conserved. In the Appendix we show what happens when the box is isolated, given mass, and allowed to recoil. Second, if the center of mass is absent, the only degree of freedom that could carry the recoil resulting from an elastic interior electron deflection is also absent. Phonons will be forced to take on that role and indeed by construction there can be no elastic events of this sort.
If the crystal total mechanical momentum is retained, and total momentum conservation is again enforceable, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. Those degrees of freedom include the several lattice zero frequency recoil modes, and internal vibrational modes.
An instructive analog is provided by Mössbauer scattering, in which the dominant elastic line, perhaps 97% of the total probability, implies recoil of the entire lattice without internal phonon excitation.
We hasten to point out that the question here involves real physics, and not mere representation: are phonons created or destroyed when electrons deflect in a pure crystal, or not? That is an essential, physical difference, with profound experimental consequences.